# Probing Chiral-Symmetric Higher-Order Topological Insulators with Multipole Winding Number

### 08 Jan 2024 v1

@@ Abstract
The interplay between crystalline symmetry and band topology gives rise to unprecedented lower-dimensional boundary states in higher-order topological insulators (HOTIs). However, the measurement of the topological invariants of HOTIs remains a significant challenge. Here, we propose the multipole winding number (MWN) for chiral-symmetric HOTIs, achieved by applying a corner twisted boundary condition. The MWN, arising from both bulk and boundary states, may accurately capture the bulk-corner correspondence in finite systems. To address the measurement challenge, we leverage the perturbative nature of the corner twisted boundary condition and develop a real-space approach for determining the MWN in both two-dimensional and three-dimensional systems. The real-space formula provides an experimentally viable strategy for directly probing the topology of chiral-symmetric HOTIs through dynamical evolution. Our findings not only highlight the twisted boundary condition as a powerful tool for investigating HOTIs, but also establish a paradigm for exploring real-space formulas for the topological invariants of HOTIs.
@@
Introduction. In contrast to conventional topological insulators, higher-order topological insulators (HOTIs) [1, 2, 3] are defined by higher-order topological invariants, resulting in the presence of lower-dimensional boundary states. The emergence of these boundary states is typically associated with non-trivial higher-order bulk topology [4]. Recent experiments have successfully demonstrated the realization of HOTIs [5, 6, 7, 8, 9, 10, 11], as evidenced by the appearance of corner states. Up to now, measuring the topological invariants of HOTIs remains a significant challenge.
@@
In addition to exploring edge topology, investigating bulk topology is fundamental. Utilizing the real-space formula [12, 13], the winding number for one-dimensional (1D) chiral-symmetric topological insulators is obtained by measuring mean chiral displacement [14, 15, 16, 17, 18, 19, 20, 21, 22]. Recently, a novel approach proposed capturing the topology of chiral-symmetric HOTIs introduces a real-space multipole chiral number based upon the difference in the multipole moments of two sublattices [23]. Chiral symmetry has been demonstrated to safeguard non-trivial higher-order topology independently of crystalline symmetry [24, 25, 23, 26]. The multipole chiral number represents a generalization of the winding number for conventional topological insulators and effectively characterizes the topology in real space. There is a suggestion to extend the method of measuring mean chiral displacement for 1D systems to 2D chiral-symmetric HOTIs by combining winding numbers in x and y directions [27]. However, this method remains within the realm of conventional topological insulators. To date, no established method exists for directly probing the topology of chiral-symmetric HOTIs from their bulk. The question persists: Can we define a winding number for chiral-symmetric HOTIs?
@@
In this Letter, we present a comprehensive and experimentally viable framework for probing chiral-symmetric HOTIs using a real-space topological invariant. We define the multipole winding number (MWN) based on the bulk-corner correspondence, a concept lacking a momentum-space counterpart. This topological invariant is defined through the corner twisted boundary condition (CTBC), linking corners and addressing gauge fields during tunneling across different corners. The MWN precisely quantifies the number of zero-energy modes in both 2D and 3D chiral-symmetric HOTIs under open boundary conditions. We introduce appropriate twist operators that transform the gauge field at corners into the bulk, rendering the gauge field a reliable perturbative parameter. Similar to deriving real-space winding numbers for 1D chiral-symmetric topological insulators [28] and real-space Chern numbers for 2D topological insulators [29], we develop a real-space formula for the MWN with perturbative expansion. We discover that the MWN is influenced not only by the bulk but also by the edge. Additionally, we demonstrate numerically how to extract the MWN through dynamical evolution. Our proposed scheme can be easily implemented using currently available experimental techniques.
Corner twisted boundary condition and multipole winding number. Usually, a chiral-symmetric lattice system can be divided into two sublattices obeying a tight-binding Hamiltonian: $\hat{H}={{\bm{\psi}}_{A}}h{\bm{\psi}}_{B}^{\dagger}+{{\bm{\psi}}_{B}}{h^{\dagger}}{\bm{\psi}}_{A}^{\dagger}$. Here, ${{\bm{\psi}}_{\alpha}}\equiv\left({|{1_{\alpha}}\rangle,|{2_{\alpha}}\rangle,\cdots}\right)$ with $\alpha=\{A,B\}$ indexing the two sublattices and $|i_{\alpha}\rangle$ being the single-particle position-space basis for the $\alpha$ sublattice. For convenience, we assume the dimensions of $A$ and $B$ sublattices are equal. The Hamiltonian is completely determined by the off-diagonal matrix $h$. The chiral symmetry results in a symmetric spectrum: the eigenvalues of $H$ appear pairwise with opposite values, which are related to the singular values of $h$ [28]. The eigen-equation $\hat{H}|{\Psi_{n}}\rangle={E_{n}}|{\Psi_{n}}\rangle$ can be expressed as two coupled equations: $hu_{B}^{[n]}={E_{n}}u_{A}^{[n]}$ and ${h^{\dagger}}u_{A}^{[n]}={E_{n}}u_{B}^{[n]}$. Here, the $n$th eigenstate reads as $|{\Psi_{n}}\rangle=\frac{1}{{\sqrt{2}}}\sum\nolimits_{j}{({u_{A}^{j,n}|{j_{A}}\rangle+u_{B}^{j,n}|{j_{B}}\rangle})}$, and the column vectors $u_{A,B}^{[n]}$ have $u^{j,n}_{A,B}$ as their $j$th elements. By performing the singular value decomposition: $h={u_{A}}\Sigma{u_{B}}^{-1}$, where $\Sigma=\mathrm{diag}(E_{1},\cdots,E_{n},\cdots)$ is diagonal-positive-definite and $u_{A,B}=[u_{A,B}^{[1]},u_{A,B}^{[2]},\cdots,u_{A,B}^{[L/2]}]$ is an unitary matrix, one can solve the coupled equations.
@@
Chiral-symmetric topological insulators in odd dimensions are characterized by the winding number [30]. In 1D, the winding number can be regarded as the dipole polarization between the two sublattices [28]. With a non-trivial winding number in bulk, zero-energy edge states emerge under OBC, which is called the bulk-edge correspondence. To capture the dipole quantity of the bulk, one can introduce a general twisted boundary condition (TBC) [31, 28]. By introducing a parameter $r$ on the tunneling across the boundary [31], which connects the OBC and the TBC, the winding number $\nu$ is determined by the closed line integral of $z\equiv\det h(r,\theta)\in\mathbb{C}$,
@@
which requires $h$ to be invertible (i.e. all its singular values are non-zero) on the path. The winding number $\nu$ reflects the fold of zero points $(z=0)$ encircled by the curve $\mathcal{C}$ on the complex plane. Meanwhile, since $\det h(r,\theta)\propto\prod\nolimits_{n}{{E_{n}}}$, these zero points are related to the number of zero singular values, which is half of the number of zero-energy modes $N_{E=0}$, i.e. $N_{E=0}=2\nu$. Thus, the winding number defined through TBC can well characterize the number of zero-energy edge states.
@@
Figure 1: (a-b) Schematic for corner twisted boundary condition. In 2D square lattices, red arrows indicate the tunneling from one corner to the other, which are parametrized by $r\in[0,1]$ and $\theta\in[0,2\pi]$. In 3D cubic lattices, the CTBC is imposed on each face, with the dashed line indicating the orientation of the twist angle. (c-d) 2D and 3D Benalcazar-Bernevig-Hughes model under OBC. The tunneling amplitudes $\gamma$ and $\lambda$ are marked by different colors. The dashed lines represent negative amplitudes.
@@
To meet the aforementioned criteria, we consider a corner twisted boundary condition (CTBC) shown in Fig. 1 (a-b), in which particles are allowed to tunnel between corners with a complex amplitude. In 2D square lattices, we adopt the CTBC [32]. In 3D cubic lattices, we consider the corner tunneling in each face of the cube, in which the orientation of the CTBC is implied by the dashed line. The CTBC is equivalent to threading a gauge field, resembling the general TBC. This choice of gauge field ensure that it will not significantly change the spectrum, which is confirmed by perturbative analysis later. To manifest the bulk-corner correspondence, the tunneling strength between corners is required to multiply a factor $r\in[0,1]$. With these conditions, we can define the MWN via Eq. (1). In addition, the CTBC can be used to define the multipole polarization [32], which is quantized by the chiral symmetry [24, 25] (see details in Supplemental Material [33] ).
Benalcazar-Bernevig-Hughes model. To illustrate the validity of our MWN, we consider the Benalcazar-Bernevig-Hughes (BBH) model [34] (see Fig. 1 (c-d)). This model preserves chiral symmetry and exhibits zero-energy corner states from higher-order topology. Notably, the BBH model with long-range tunneling has been realized in experiments [26]. In 2D, the system obeys the tight-binding Hamiltonian: ${{\hat{H}}_{{\rm{2D}}}}=-\sum\nolimits_{x,y}{\left[{{t_{x}}\hat{a}_{x+1,y}^{\dagger}{{\hat{a}}_{x,y}}+{{\left({-1}\right)}^{x}}{t_{y}}\hat{a}_{x,y+1}^{\dagger}{{\hat{a}}_{x,y}}}\right]}+{\rm{H}}.{\rm{c}}$, with ${t_{x}}=\frac{{\gamma+\lambda}}{2}+\frac{{\gamma-\lambda}}{2}{\left({-1}\right)^{x}},{t_{y}}=\frac{{\gamma+\lambda}}{2}+\frac{{\gamma-\lambda}}{2}{\left({-1}\right)^{y}}$. It is known that there are two topological phases determined by the ratio $|\gamma/\lambda|$. If $|\gamma/\lambda|<1$, a non-zero bulk multipole polarization arises and zero-energy modes emerge at corners under the OBC. Otherwise, if $|\gamma/\lambda|>1$, the system is well gapped and there is no in-gap zero modes. From the multipole polarization under periodic boundary condition (PBC), the phase boundary can be given as $|\gamma/\lambda|=1$. Under the CTBC, we calculate the spectrum and the MWN, see Fig. 2. Interestingly, the phase boundary deviates from $\gamma/\lambda=1$ due to the finite-size effect. Instead, the MWN correctly captures the gap closing point under the CTBC. The localization length of corner states near the phase transition point becomes significantly large, leading to hybridization of corner states and resulting in a non-zero energy shift. Thus, the phase transition point given by the MWN is also shifted. However, in the thermodynamic limit $L\to\infty$, the phase transition point approaches to $\gamma/\lambda=1$, see the inset in Fig. 2 (b). This phenomenon can by accounted by the topological obstruction between PBC and OBC [35].
@@
Subsequently, we calculate the lowest singular value $\min({\Sigma})$ and the determinant $\det h(r,\theta)$ as functions of $r$ and $\theta$, see the right panel of Fig. 2. The lowest singular value of $h(r,\theta)$ corresponds to the eigenenergy most close to $E=0$. For $|\gamma/\lambda|<1$ (and away from the phase transition point), we confirm that there appears $2$ ($4$)-fold degenerate singular points at $r=0$ in the 2D (3D) BBH model. This is precisely captured by the MWN. Near the phase transition point, the singular point gradually move out of the integral circle, causing a zero MWN.
@@
Figure 2: (a) Spectrum and (b) MWN of the 2D BBH model versus $\gamma/\lambda$ under the CTBC. Inset in (b): the critical $\gamma/\lambda$ versus $1/L$ with a red dashed line for guidance. Right panel: the lowest singular values $\min(\Sigma)$ of $h(r,\theta)$ on the complex plane for $L=20$. Arrows represent the angle of $\det h(r,\theta)$.
@@
Real-space representation of multipole winding number. Below we show how to derive the real-space formulation for the MWN. The central idea is to make use of the gauge freedom of the CTBC and find appropriate gauges that makes the twist angle distributed uniformly in the bulk. Then the twist angle can be treated as a perturbation in order of $\theta/L$ and one can use the perturbative expansion to derive the MWN. To perform the gauge transformation, we introduce the twist operators
@@
for 3D systems. Through the transformation $\tilde{H}\left(\theta\right)=\hat{U}_{\theta}\hat{H}\left(\theta\right){{\hat{U}_{\theta}}^{-1}}$, the gauge fields at corners are transformed into the bulk with diluted and uniform strength (see details in Supplemental Material [33]). The Hamiltonian $\hat{H}\left(\theta\right)$ under the CTBC satisfies the gauge-invariant relation: $\hat{H}\left({\theta+2\pi}\right)=\hat{H}\left(\theta\right)$, while the transformed Hamiltonian satisfies the covariant relation: $\tilde{H}\left({\theta+2\pi}\right)={{\hat{U}}_{2\pi}}\tilde{H}\left(\theta\right){{\hat{U}}_{2\pi}}^{-1}$. For simplicity, we refer to the former one as “corner gauge” and to the latter as “bulk gauge”. Operators under the bulk gauge will be denoted by tilde sign: $h\to\tilde{h}$. Meanwhile, the twist angle (gauge field) will only cause a slight modification in the spectrum, which can be neglected in the thermodynamic limit.
Choosing $r=1$, under the bulk gauge, the MWN can still be expressed as Eq. (1). Keeping the terms up to the first order of $1/L$, we have $\tilde{h}\left(\theta\right)=h\left(0\right)+\frac{\theta}{L}{[{\frac{\partial}{{\partial\theta/L}}\tilde{h}\left(\theta\right)}]_{\theta=0}}+O\left({\frac{1}{{{L^{2}}}}}\right)$ and ${\rm{Tr}}[{h{{\left({r,\theta}\right)}^{-1}}{\partial_{\theta}}h\left({r,\theta}\right)}]\approx{\rm{Tr}}\{{{h^{-1}}\left(0\right){{[{{{\partial_{\theta}}}\tilde{h}\left(\theta\right)}]}_{\theta=0}}}\}$ being a constant. Therefore, the MWN can be given as $\nu\approx-i{\rm{Tr}}\{{{h^{-1}}\left(0\right){{[{{{\partial_{\theta}}}\tilde{h}\left(\theta\right)}]}_{\theta=0}}}\}$. Using $\tilde{h}\left(\theta\right)={\bm{\psi}}_{A}^{\dagger}\tilde{H}\left(\theta\right){{\bm{\psi}}_{B}}$, we have $\tilde{h}={\cal U}_{2\pi}^{A}\tilde{h}\left(0\right){({{\cal U}_{2\pi}^{B}})^{-1}}$ with ${{\hat{U}}_{2\pi}}=\exp({i2\pi\hat{\mathcal{P}}/L})$, where $\hat{\mathcal{P}}=\{\hat{\cal{P}}_{xy},\hat{\cal{P}}_{xyz}\}$ are both diagonal in the position basis. For simplicity, we define the projected twist operator matrices in the sector $\alpha=\{A,B\}$: ${\cal U}_{2\pi}^{\alpha}\equiv{\bm{\psi}}_{\alpha}^{\dagger}{{\hat{U}}_{2\pi}}{{\bm{\psi}}_{\alpha}}$. To ensure the gauge invariance, we impose the constraint $\det\tilde{h}\left({2\pi}\right)=\det\tilde{h}\left(0\right)$. Thus, we require ${\cal U}_{2\pi}\equiv{\cal U}_{2\pi}^{A}={\cal U}_{2\pi}^{B}$, which can be easily satisfied by setting an identical position for all sublattices within a unit cell. Up to the first order of $1/L$, we have ${\left[{{\partial_{\theta}}\tilde{h}\left(\theta\right)}\right]_{\theta=0}}\approx\frac{{\tilde{h}\left({2\pi}\right)-\tilde{h}\left(0\right)}}{{2\pi}}=\frac{{{\cal U}_{2\pi}h{{{{\cal U}_{2\pi}^{-1}}}}-h}}{{2\pi}}$. To gain higher accuracy, one may consider a higher-order finite difference (see details in Supplemental Material [33]). Then, a rather simpler real-space formula can be obtained: $\nu=\frac{1}{{2\pi i}}{\rm{Tr}}\left({{h^{-1}}{\cal U}_{2\pi}h{{{{\cal U}_{2\pi}^{-1}}}}-I}\right)$. This leads to two equivalent real-space formulas: (i) the Bott index form and (ii) the non-commutative form.
Since ${{h^{-1}}{\cal U}_{2\pi}h{{{{\cal U}_{2\pi}^{-1}}}}}$ is close to an identity matrix, in the thermodynamic limit $L\to\infty$, one can obtain the Bott index form $\nu=\frac{1}{{2\pi i}}{\rm{Tr}}\left[{\log\left({{h^{-1}}{\cal U}_{2\pi}h{{{{\cal U}_{2\pi}^{-1}}}}}\right)}\right]$. This formula is analogous to the multipole chiral number [23], which also takes the Bott index form. The multipole chiral number is defined based on the quadrupole and octupole of polarization. It can be proved that the multipole chiral number is equivalent to the real-space winding number defined through a modified CTBC. The modified CTBC still connects the corners, while only those tunnelings related to one of the corners are threaded by a twist angle. In this scenario, the modified MWN captures the number of zero-energy modes induced in one specific corner (see details in Supplemental Material [33]). Therefore, although the generators of the twist operators [Eq. (2) and Eq. (Probing Chiral-Symmetric Higher-Order Topological Insulators with Multipole Winding Number)] are quite different from the quadrupole and octupole, they share the same physical origin.
Using the Baker-Campbell-Hausdorff (BCH) formula, we have ${{\cal U}_{2\pi}}h{{\cal U}_{2\pi}^{-1}}\approx h+2\pi i\left[{P,h}\right]/L$ with $P={\bm{\psi}}_{A}^{\dagger}\hat{\mathcal{P}}{\bm{\psi}}_{A}$ denoting the generator projected onto the sector $A$, and then obtain the non-commutative form $\nu=\frac{1}{L}\mathrm{Tr}\left({h^{-1}}\left[{{{P}},h}\right]\right)$. However, this approximation only holds for systems with OBC and $L\to\infty$. In a finite system with CTBC, this approximation is poor because it always evaluates to zero if $h$ is invertible: ${\rm{Tr}}\left({{h^{-1}}\left[{P,h}\right]}\right)={\rm{Tr}}\left({{h^{-1}}Ph-P}\right)=0$. This issue arises from the slow convergence of the BCH formula and the omission of higher-order terms (see details in Supplemental Material [33]). Meanwhile, as the CTBC connects the corners, general definition of the position near the corner become ill-defined, which also leads to the poor performance of the first-order expansion. However, under the OBC, the BCH formula converges fast because position-dependent operators are well-defined. In the presence of zero modes, $h$ becomes singular. In this case, we need to rule out the zero singular values and the associated vectors and use the pseudo-inverse formulation: $h^{-1}\approx\tilde{u}_{B}\tilde{\Sigma}^{-1}\tilde{u}_{A}^{{\dagger}}$ and then the non-commutative form will be non-zero.
Measurement of multipole winding number. The above forms for MWN are not straightforward to be measured in experiments. Below we derive a formula suitable for experimental measurement. Firstly, the flattened Hamiltonian can be written as $\hat{Q}={{\hat{\mathbb{P}}}_{+}}-{{\hat{\mathbb{P}}}_{-}}$ with ^${{\hat{\mathbb{P}}}_{-}}=\sum\nolimits_{{E_{n}}<0}{|{\psi_{n}}\rangle\langle{\psi_{n}}|}$ and ^${{\hat{\mathbb{P}}}_{+}}=\sum\nolimits_{{E_{n}}>0}{|{\psi_{n}}\rangle\langle{\psi_{n}}|}$, where $E_{n}$ is the eigenenergy for the eigenstate $|\psi_{n}\rangle$. The flattened Hamiltonian shares the same eigenstates of the original Hamiltonian and the zero-energy mode is naturally excluded. Replacing the off-diagonal matrix $h$ with the flattened matrix $q\equiv{\bm{\psi}}_{A}^{\dagger}\hat{Q}{{\bm{\psi}}_{B}}$ and $q^{-1}\equiv{\bm{\psi}}_{B}^{\dagger}\hat{Q}{{\bm{\psi}}_{A}}$, we have $\nu=\frac{1}{L}\mathrm{Tr}\left({q^{-1}}\left[{{{P}},q}\right]\right)$. Therefore, we obtain a real-space formula (see details in Supplemental Material [33]),

### 21 Mar 2024 v2

@@ Abstract
The interplay between crystalline symmetry and band topology gives rise to unprecedented lower-dimensional boundary states in higher-order topological insulators (HOTIs). However, the measurement of the topological invariants of HOTIs remains a significant challenge. Here, we define a multipole winding number (MWN) for chiral-symmetric HOTIs by applying a corner twisted boundary condition. The MWN, arising from both bulk and boundary states, accurately captures the bulk-corner correspondence including boundary-obstructed topological phases. To address the measurement challenge, we leverage the perturbative nature of the corner twisted boundary condition and develop a real-space approach for determining the MWN in both two-dimensional and three-dimensional systems. The real-space formula provides an experimentally viable strategy for directly probing the topology of chiral-symmetric HOTIs through dynamical evolution. Our findings not only highlight the twisted boundary condition as a powerful tool for investigating HOTIs, but also establish a paradigm for exploring real-space formulas for the topological invariants of HOTIs.
Introduction
@@
In contrast to conventional topological insulators, higher-order topological insulators (HOTIs) [1, 2, 3] are characterized by higher-order topological invariants, resulting in the presence of lower-dimensional boundary states. The emergence of these boundary states is typically associated with non-trivial higher-order bulk topology [4]. Recent experiments have successfully demonstrated the realization of HOTIs [5, 6, 7, 8, 9, 10, 11, 12], as evidenced by the appearance of corner states. However, the measurement of topological invariants of HOTIs still remains a significant challenge, as topological invariants are a global quantity, which are usually related to the bulk.
@@
In addition to exploring the edge topology, it is crucial to characterize bulk topology. In one-dimensional (1D) chiral-symmetric topological insulators, it is known that the winding number can be obtained by measuring mean chiral displacement in the bulk [13, 14, 15, 16, 17, 18, 19]. This approach is based on the real-space formula of the winding number [20, 21]. Recently, the multipole chiral number (MCN) approach is proposed to capture the topology of chiral-symmetric HOTIs based upon the difference in the multipole moments of two sublattices in real space [22]. Chiral symmetry has been demonstrated to safeguard non-trivial higher-order topology independently of crystalline symmetry [23, 24, 22, 25, 26]. The MCN represents a generalization of the real-space winding number for conventional topological insulators [27] and effectively characterizes the higher-order topology. Remarkably, the MCN is a $\mathbb{Z}$ number for characterizing chiral-symmetric topological insulators. This is in stack contrast to the multipole polarization, which only gives rise to a $\mathbb{Z}_{2}$ quantity.
Up to now, the experimental investigation of the chiral-symmetric HOTIs with larger topological numbers relies on identifying the corner states [25, 28, 29]. To probe the bulk topology of HOTIs, it has been proposed to generalize the concept of mean chiral displacement. By combining the winding numbers in $x$ and $y$ directions [30], the method of measuring mean chiral displacement for 1D systems can be extended to 2D chiral-symmetric HOTIs. However, this method remains within the realm of conventional topological insulators. Besides, a generalization of the mean chiral displacement to the quadrupole insulator [31] successfully extracts the quadrupole nature of the chiral-symmetric HOTI from the bulk. Given the similarities between the first-order chiral-symmetric topological insulator and the higher-order HOTI, a question arises: Can we define a winding number for chiral-symmetric HOTIs? Furthermore, can we directly measure such a winding number?
@@
In this article, we present a comprehensive and experimentally viable framework for probing chiral-symmetric HOTIs using a real-space topological invariant. We define the multipole winding number (MWN), a concept lacking a momentum-space counterpart, based on the bulk-corner correspondence. This topological invariant is defined through the corner twisted boundary condition (CTBC), linking corners and addressing gauge fields during tunneling across different corners. It precisely quantifies the number of zero-energy modes in both 2D and 3D chiral-symmetric HOTIs under open boundary conditions even for the boundary-obstructed topological phase (BOTP). Remarkably, the MWN and the corner modes are found to be contributed from both bulk and boundary states. Similar to deriving real-space winding numbers for 1D chiral-symmetric topological insulators [27] and real-space Chern numbers for 2D topological insulators [32], we develop a real-space formula for the MWN via perturbative expansion. We elaborate appropriate twist operators that transform the gauge field at corners into the bulk, rendering the gauge field a reliable perturbative parameter. In further, we demonstrate numerically how to extract the MWN through dynamical evolution. Our proposed scheme can be easily implemented using currently available experimental techniques.
Results
@@
Corner twisted boundary condition and multipole winding number. It is known that chiral-symmetric topological insulators in odd dimensions are characterized by the winding number [33]. In 1D, the winding number can be regarded as the dipole polarization between the two sublattices [27]. With a non-trivial winding number in bulk, zero-energy edge states emerge under OBC, which is called the bulk-edge correspondence. To capture the dipole quantity of the bulk, one can introduce a general twisted boundary condition (TBC) [34, 27]. By introducing a parameter $r$ on the tunneling across the boundary [34], which connects the OBC and the TBC, the winding number $\nu$ is determined by the closed line integral of $z\equiv\det h(r,\theta)\in\mathbb{C}$,
@@
which requires $h$ to be invertible (i.e. all its singular values are non-zero) on the path. The winding number $\nu$ reflects the fold of zero points $(z=0)$ encircled by the curve $\mathcal{C}$ on the complex plane. Meanwhile, since $\det h(r,\theta)\propto\prod\nolimits_{n}{{E_{n}}}$ (see Methods for proof), these zero points are related to the number of zero singular values, which is half of the number of zero-energy modes $N_{E=0}$, i.e. $N_{E=0}=2\nu$. Thus, the winding number defined through TBC can well characterize the number of zero-energy edge states.
@@
Figure 1: Schematic for corner twisted boundary condition and the Benalcazar-Bernevig-Hughes model. a 2D case. The red arrows indicate the tunneling from one corner to the other, which are parametrized by $r\in[0,1]$ and $\theta\in[0,2\pi]$. b 3D case. The CTBC is imposed on each face, with the dashed line indicating the orientation of the twist angle. c-d 2D and 3D Benalcazar-Bernevig-Hughes model under OBC. The tunneling amplitudes $\gamma$ and $\lambda$ are marked by different colors. The dashed lines represent negative amplitudes.
@@
To meet the aforementioned criteria, we consider a corner twisted boundary condition (CTBC) shown in Fig. 1 (a-b), in which particles are allowed to tunnel between corners with a complex amplitude. In 2D square lattices, we adopt the CTBC [35]. In 3D cubic lattices, we consider the corner tunneling in each face of the cube, in which the orientation of the CTBC is implied by the dashed line. The CTBC is equivalent to threading a gauge field, resembling the general TBC. This choice of gauge field ensure that it will not significantly change the spectrum, which is confirmed by perturbative analysis later. To manifest the bulk-corner correspondence, the tunneling strength between corners is required to multiply a factor $r\in[0,1]$. With these conditions, we can define the MWN via Eq. (1). In addition, the CTBC can be used to define the multipole polarization [35], which is quantized by the chiral symmetry [23, 24] (see proof in Sec. A of Supplementary Note 1.).
Benalcazar-Bernevig-Hughes model. To illustrate the validity of our MWN, we consider the Benalcazar-Bernevig-Hughes (BBH) model [36] (see Fig. 1 (c-d)). This model preserves chiral symmetry and exhibits zero-energy corner states from higher-order topology. Notably, the BBH model with long-range tunneling has been realized in experiments [25, 28, 29]. In 2D, the system obeys the tight-binding Hamiltonian:
${{\hat{H}}_{{\rm{2D}}}}=-\sum\limits_{x,y}{\left[{{t_{x}}\hat{a}_{x+1,y}^{\dagger}{{\hat{a}}_{x,y}}+{{\left({-1}\right)}^{x}}{t_{y}}\hat{a}_{x,y+1}^{\dagger}{{\hat{a}}_{x,y}}}\right]}+{\rm{H.c.}},$ (2)
with ${t_{x}}=\frac{{\gamma+\lambda}}{2}+\frac{{\gamma-\lambda}}{2}{\left({-1}\right)^{x}},{t_{y}}=\frac{{\gamma+\lambda}}{2}+\frac{{\gamma-\lambda}}{2}{\left({-1}\right)^{y}}$. It is known that there are two topological phases determined by the ratio $|\gamma/\lambda|$. If $|\gamma/\lambda|<1$, a non-zero bulk multipole polarization arises and zero-energy modes emerge at corners under the OBC. Otherwise, if $|\gamma/\lambda|>1$, the system is well gapped and there is no in-gap zero modes. From the multipole polarization under periodic boundary condition (PBC), the phase transition point can be given as $|\gamma/\lambda|=1$.
Under the CTBC, we calculate the spectrum and the MWN, see Fig. 2 (a-b). Interestingly, the phase transition point, which is characterized by the gap closing and the sudden change of MWN, deviates from $\gamma/\lambda=1$. This can be attributed to the finite-size effect. Under OBC, the localization length of corner states near the phase transition point becomes significantly large, leading to hybridization of corner states and resulting in a non-zero energy shift in finite systems. According to Eq. (1), the MWN is related to the number of zero-energy mode under OBC inside the integral circle ($r=1$). Near the phase transition point, the singular point gradually move out of the integral circle due to the energy shift, causing a change of MWN. Thus, the phase transition point given by the MWN is also shifted due to the finite-size effect (see detailed discussions in Sec. B of Supplementary Note 2). In the thermodynamic limit $L\to\infty$, the phase transition point approaches to $\gamma/\lambda=1$, as shown in the inset of Fig. 2 (b). This shows that the MWN can well capture the bulk topology.
@@
Subsequently, we calculate the lowest non-negative energy $\min({|E|})$ and the determinant $\det h(r,\theta)$ as functions of $r$ and $\theta$, see the right panel of Fig. 2. For $|\gamma/\lambda|<1$ (and away from the phase transition point), we confirm that there appears $2$ ($4$)-fold degenerate singular points at $r=0$ in the 2D (3D) BBH model. These singular points can be viewed as topological defect, which are precisely captured by the MWN and consistent with half the number of corner states under OBC.
@@
Figure 2: Energy spectrum and the MWN of the 2D BBH model and schematic illustration of the winding number. a Energy Spectrum computed from exact diagonalization as a function of $\gamma/\lambda$ under the CTBC. b The corresponding MWN. Inset in b: the critical $\gamma/\lambda$ versus $1/L$ with a red dashed line for guidance. c The lowest singular values $\min(\Sigma)$ of $h(r,\theta)$ on the complex plane for $L_{x}=L_{y}=L=20$. Arrows represent the angle of $\det h(r,\theta)$.
Boundary-obstructed topological phase. Furthermore, we demonstrate that the MWN can effectively characterize the topology and the bulk-corner correspondence for boundary-obstructed topological phase (BOTP). It has been shown that different BOTPs can be deformed continuously to each other without closing the bulk gap under the PBC [37]. Nevertheless, the phase transition between different BOTPs is associated with gap closing under the OBC. To realize a BOTP, we break the $C_{4}$ symmetry in the 2D BBH model by letting ${t_{x}}=\frac{{\gamma_{x}+\lambda}}{2}+\frac{{\gamma_{x}-\lambda}}{2}{\left({-1}\right)^{x}},{t_{y}}=\frac{{\gamma_{y}+\lambda}}{2}+\frac{{\gamma_{y}-\lambda}}{2}{\left({-1}\right)^{y}}$ in Eq. (2). In this model, zero-energy modes emerge under OBC when $|\gamma_{x,y}/\lambda|<1$. If either $|\gamma_{x}/\lambda|>1$ or $|\gamma_{y}/\lambda|>1$, no zero-energy mode exists, indicating that $|\gamma_{x,y}/\lambda|=1$ are phase boundaries. However, under PBC, the spectral gap only closes at some points where $|\gamma_{x}/\lambda|=|\gamma_{y}/\lambda|=1$. Thus, the phase transition between BOTPs can not be characterized by the gap closing of the bulk. Instead, by employing CTBC, we observe a distinct gap-closing point nearer either $|\gamma_{x}/\lambda|=1$ or $|\gamma_{y}/\lambda|=1$, as shown in Fig. 3. Remarkably, the MWN accurately captures the presence of zero-energy modes and the phase transition of BOTP under OBC. This is because the gap closing is resulted not only from the bulk states but also the boundary states under CTBC. The non-trivial MWN is contributed from both of them when $|\gamma_{x,y}|<1$ Therefore, it can be inferred that the corner state is resulted from the non-trivial topology of both the bulk and boundary, which will be examined in real space below.
Figure 3: Energy spectrum and the MWN of the 2D BBH model under the boundary obstructed topological phase (BOTP). a-c Energy Spectrum computed from exact diagonalization under the PBC, CTBC, and OBC respectively. d The corresponding MWN computed via Eq. (1) under CTBC. Other parameters are set to: $\gamma_{y}/\lambda=0.5$ and $L=40$.
Real-space representation of multipole winding number. In experiments, the simplest method to distinguish chiral-symmetric HOTIs from trivial topological insulators is through observing their corner states. In addition to exploring corner states resulted from non-trivial bulk topology, researchers are also making efforts to extract the topological invariant of HOTIs experimentally [38, 30, 39, 31]. Below, we derive a real-space formulation for MWN and explore feasible method for corresponding experimental measurement.
@@
To attain the real-space formula, our central idea is to make use of the gauge freedom of the CTBC and find appropriate gauges that makes the twist angle distributed uniformly in the bulk [27, 32]. Then the twist angle can be treated as a perturbation in order of $\theta/L$ and one can use the perturbative expansion to derive the MWN. To perform the desired gauge transformation, we introduce the twist operators
@@
for 3D systems. Through the transformation $\tilde{H}\left(\theta\right)=\hat{U}_{\theta}\hat{H}\left(\theta\right){{\hat{U}_{\theta}}^{-1}}$, the gauge fields at corners are transformed into the bulk with diluted and uniform strength (see illustrations in Sec. C of Supplementary Note 1). The Hamiltonian $\hat{H}\left(\theta\right)$ under the CTBC satisfies the gauge-invariant relation: $\hat{H}\left({\theta+2\pi}\right)=\hat{H}\left(\theta\right)$, while the transformed Hamiltonian satisfies the covariant relation: $\tilde{H}\left({\theta+2\pi}\right)={{\hat{U}}_{2\pi}}\tilde{H}\left(\theta\right){{\hat{U}}_{2\pi}}^{-1}$. For simplicity, we refer to the former one as “corner gauge” and to the latter as “bulk gauge”. Operators under the bulk gauge will be denoted by tilde sign: $h\to\tilde{h}$. By keeping up to the first order of $1/L$ in Eq. (1), we finally obtain the following real-space expression for the MWN (see detailed derivation in Methods):